1-

If X, . . . , X22 is a random sample from a normal distribution with parameters μ = 18.4 and ơa_ 13.3, find P(164.11-Σ(x-18.4)2-535.86)

2-
Answer #1

Question 1 The given question represents the chi-square distibution in definition. Here,

the value of \sum_{i=1}^{22}(X_i - 18.4)^2 = Zi2 * σ2

so as we know that

X2i =  \sum_{i=1}^{k}Z_i^2

so here k = 22 degree of freedom and we can rewrite the given probability as

P(164.11/13.3 < \sum_{i=1}^{22}Z_i^2 < 535.86/13.3) = P(12.339 < \sum_{i=1}^{k}Z_i^2 < 40.365)

= 0.95 - 0.01 = 0.94

as these probability can be calculated by CHIDIST formula or chi - square table .

(2) Here Pr(right handed) = 0.80

Pr(Left handed) = 0.16

Pr(Ambidextrous) = 0.04

Pr(121 RH, 22 LH , 7 ambi) = 150C121 (0.80)121* 29C22 (0.16)22 * 7C7 (0.04)7 = 0.01187